Assume that the market for Good X is defined as follows: QD = 64 - 16P and QS = 16P - 8. If the government imposes a price floor in this market at $3.00, what will producer surplus be?
Answer #1

Put D= 3 is 0= 64. 162 64-1613) when Q=16, 2=180-8 16+8=161 P-1.5 tradnes Surplus = (3-1.5) 16 + 1 (1.5-0.5)416 24 +8 = 32 32